OpenStudy (anonymous):

the perimeter of a rectangle is 28 cm. the length of the rectangle is 2 cm more than twice the width. find the dimension of the rectangle

OpenStudy (anonymous):

So start off with the equation Perimeter= 2(length)+2(width). Plug in numbers/variables, and lets use the letter "w" to represent width: 28=2(2w+2)+2w Divide the entire equation by 2: 14=2w+2+w 14=3w+2 12=3w w=4 Now thats only your width, so now find your length: 2(4)+2=10 So your length is 10, and width is 4. you can even check: 2(10)+2(4)=20+8=28